If x is an odd whole number, which of the following also represents an odd number?
a. 2x+1
b. x-2
c. 4x-3
d. all of the above

Answer by Fombitz(32388)

a.)Twice an odd number gives you an even number.
If I add 1 to an even number I get an odd number.
Example : 2(3)+1 = 7
b.)2 subtract(or added) to an odd number gives another odd number.
Example : 5-2=3
c.)Four times an odd number gives you an even number.
Subtracting 2 would give you an even number.
Subtracting 3 would give you an odd number.
Example : 4(11)-3=44-3=41.
The answer is d.) All of them.